OpenStudy (anonymous):

When graphing the following equations, which one will be a straight line? A. y = x^2 + 2 B. f(x) =2/3 x + 1 C. y = |x + 4| D. x^2 + y^2 = 1

OpenStudy (anonymous):

A straight line is a function of the form y = mx+b Which choice is correct?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

I said y = mx + b, where m is the slope and b = y-intercept.

OpenStudy (anonymous):

Which is of the form y = mx+b?

OpenStudy (anonymous):

a is

OpenStudy (anonymous):

How is x^2 of that form?????????????????????

OpenStudy (anonymous):

I said y = mx + b, NOT x^2

OpenStudy (anonymous):

Choice B is of that form..it is identical to y = mx + b.

OpenStudy (anonymous):

k